# ConsDreamer: Advancing Multi-View Consistency for Zero-Shot Text-to-3D Generation

### v1

@@ III Methodology
To tackle the multi-face Janus problem, we propose ConsDreamer, a unified framework to enhance view semantic clarity. Fig.3 provides an overview of our approach. Section III-A presents a mathematical analysis of the problem, identifying two problematic terms: the conditional and unconditional terms, both of which introduce prior view preferences during the score distillation process. Based on this analysis, Section III-B introduces VDM to eliminate prior view preferences and integrate target view control in the conditional term. Meanwhile, for the unconditional term, Section III-C explores similarity distribution patterns across views in rendered images and proposes a similarity-based partial order loss $\mathcal{L}_{P}$ to enhance spatial consistency in 3D content.
@@ III-A Analysis of the multi-face Janus problem
where $R_{n}$ is a 2D rendered image from the 3D content in the ${(n-1)}^{\text{th}}$ iteration, and $Z_{\theta,n}$ denotes the 3D representation at the $n^{\text{th}}$ iteration. In Eq.(4), the user prompt $c$ is expected only to provide content information without views information. However, the current pre-trained T2I models incorporate view prior knowledge when “understanding” user prompt $c$ even if $c$ lacks explicit view information. This prior knowledge arises from the prior preference views of training data. Therefore, we decompose $c$ into a content component $c_{c}$ and a view prior component $c_{v}$, which are expected to be orthogonal:
$\tilde{p}_{3D}(\theta)=\prod_{n=1}^{N}p\left(Z_{\theta,n}\mid\lambda_{n},c_{c},c_{v},R_{n}\right),\quad\text{s.t.}\,c_{c}\perp c_{v},$ (5)
@@
${\rm\log}\tilde{p}_{3D}(\theta)=\sum_{n=1}^{N}{\rm\log}p\left(Z_{\theta,n}\mid\lambda_{n},c_{c},c_{v},R_{n}\right),$ (6)
the gradient of $\theta$ denoted as $\nabla\theta{\rm\log}\tilde{p}_{3D}(\theta)$, can be expressed as:
$\nabla_{\theta}{\rm\log}\tilde{p}_{3D}(\theta)=\sum_{n=1}^{N}\nabla_{\theta}{\rm\log}p\left(Z_{\theta,n}\mid\lambda_{n},c_{c},c_{v},R_{n}\right)$ (7)
$=\sum_{n=1}^{N}\left(\frac{\partial{\rm\log}p\left(Z_{\theta,n}\mid\lambda_{n},c_{c},c_{v},R_{n}\right)}{\partial_{Z_{\theta,n}}}\right)\frac{\partial_{Z_{\theta,n}}}{\partial_{\theta}},$
as $c_{c}$ and $c_{v}$ are independent, Eq.(7) can be further expanded using Bayes’ theorem. Specifically, the joint distribution $p(Z_{\theta,n},\lambda_{n},c_{c},c_{v},R_{n})$ satisfies the decomposition property, allowing it to be expressed as $p(Z_{\theta,n})p(\lambda_{n},c_{c},c_{v},R_{n}\mid Z_{\theta,n})$. This ensures that the transformation using Bayes’ theorem is valid:
$\nabla_{\theta}{\rm\log}\tilde{p}_{3D}(\theta)\!=\!\sum\limits_{n=1}^{N}\!\left(\!\frac{\partial{\rm\log}p\left(\lambda_{n},c_{c},c_{v},R_{n}\!\mid\!Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}\!+\!\frac{\partial{\rm\log}p\left(Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}\right)\!\frac{\partial_{Z_{\theta,n}}}{\partial_{\theta}},\end{aligned}$
where $\frac{\partial{\rm\log}p\left(Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}$ is the unconditional score [33], [49] modelled by T2I models, intuitively reflects the model’s perception of 3D objects without external view guidance. However, as previously mentioned, such perception is prone to prior preference views, which leads to the multi-face Janus problem. To address this issue, Section III-C explores similarity distribution patterns across rendered images and proposes a similarity-based partial order loss $\mathcal{L}_{P}$ to enhance the model’s view awareness in the absence of conditional view information.
@@
$&\frac{\partial}{\partial_{z_{\theta,n}}}{\rm\log}p\left(\lambda_{n},c_{c},c_{v},R_{n}\mid Z_{\theta,n}\right)\\ &\!=\!\underbrace{\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}p\!\left(c_{c},c_{v},R_{n}\!\mid\!Z_{\theta,n}\right)}_{\text{\normalsize View Bias}}\!+\!\underbrace{\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}p\!\left(\lambda_{n},c_{c},R_{n}\!\mid\!Z_{\theta,n}\right)}_{\text{\normalsize View Control}}\!+\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}M,\end{aligned}$
where $M=\frac{p\left(\lambda_{n},c_{c},c_{v},R_{n}\mid Z_{\theta,n}\right)}{p\left(c_{c},c_{v},R_{n}\mid Z_{\theta,n}\right)p\left(\lambda_{n},c_{c},R_{n}\mid Z_{\theta,n}\right)}$.
$c_{c}$ and $R_{n}$ are constants in the current optimization step, then $M$ can be simplified by using the definition of conditional probability:
@@
Observing Eq.(9) and Eq.(10), when the view control $\lambda_{n}$ and the prior view preference $c_{v}$ conflict under the condition $Z_{\theta,n}$, i.e., $p\left(\lambda_{n}\mid c_{v},Z_{\theta,n}\right)\ll p\left(\lambda_{n}\mid Z_{\theta,n}\right)$, the view bias and the view control in Eq.(9) will have a detrimental impact on the 3D generation jointly. Additionally, in this case, $M$ approaches 0, and the third term $\nabla_{Z_{\theta,n}}{\rm\log}M$ in Eq.(9) introduces a large negative gradient, disrupting the optimization process of the model. Conversely, when $\lambda_{n}$ and $c_{v}$ are consistent, i.e., $p\left(\lambda_{n}\mid c_{v},Z_{\theta,n}\right)\approx p\left(\lambda_{n}\mid Z_{\theta,n}\right)$, the gradient directions of the first two terms in Eq.(9) align, leading to a more stable optimization process. Furthermore, as $M$ approaches 1, the gradient term $\nabla_{Z_{\theta,n}}{\rm\log}M$ approaches 0, ensuring it does not interfere with the optimization process. Therefore, establishing semantic consistency between $c_{v}$ and $\lambda_{n}$ is essential for effectively mitigating the multi-face Janus problem. Section III-B eliminates prior view preferences by disentangling view features and injects precise view control when optimizing the rendering results for specific views, thereby enhancing the semantic consistency between $c_{v}$ and $\lambda_{n}$.
@@ III-B View Disentanglement Module
Based on the mathematical analysis in Section III-A, we aim to eliminate the prior view preference and strengthen the target view specified by the camera parameters. The optimization objective can be expressed as maximizing the consistency between $c_{v}$ and $\lambda_{n}$:
@@
The process is illustrated in the upper of Fig.3. First, identify the keyword from the prompt (e.g., the keyword “squirrel” of the prompt “A squirrel eating a burger”), and feed it into the encoder to produce the view-free keyword embedding $\mathbf{V}_{{key}}^{\emptyset}$. Subsequently, new embeddings $\mathbf{V}_{{key}}^{{view}}$ corresponding to various views are obtained by combining the keyword with different view descriptions (e.g., “back view,” “side view,” and “front view”). Unlike $\mathbf{V}_{{key}}^{\phi}$, $\mathbf{V}_{{key}}^{{view}}$ incorporates auxiliary contextual information, which enables the view-specific features $\delta_{{key}}^{{view}}$ isolating content-agnostic information. The process is formulated as follows:
$\delta_{{key}}^{{view}}=\mathbf{V}_{{key}}^{{view}}-{Proj}_{\mathbf{V}_{key}^{\phi}}\left(\mathbf{V}_{{key}}^{{view}}\right),$ (12)
where $Proj_{v}(u)=\frac{u\cdot v}{||v||^{2}}v$ represents the projection of $u$ onto $v$. The view information is deprived from $\mathbf{V}_{{key}}^{{view}}$ by projecting $\mathbf{V}_{{key}}^{{view}}$ to $\mathbf{V}_{{key}}^{\emptyset}$. Furthermore, the view features $\delta_{{key}}^{{view}}$ are obtained by subtracting this projection from the $\mathbf{V}_{{key}}^{{view}}$. The $\delta_{{key}}^{{view}}$ can be applied to user prompts, enabling view control.
@@
[…]
@@
where $w_{1}$ controls the intensity of view injection, $|r|/90$ normalizes the azimuth $r$ symmetrically within the range of -90° to 90°, ensuring that the injection intensity corresponds proportionally to the deviation from the front view. When the azimuth $r$ falls within the range $(-180^{\circ},-90^{\circ})\cup(90^{\circ},180^{\circ})$, both $\delta_{{key}}^{{back}}$ and $\delta_{{key}}^{{side}}$ must be considered simultaneously, because diffusion models may be biased toward generating front views [25]:

### v2

@@ III Methodology
TABLE I: Summary of main symbols and notations used in the analysis.
Symbol Description
$c$ User-provided text prompt.
$c_{sbj}$ Subject-related component of $c$.
$c_{v}$ Prior view bias component of $c$.
$\lambda$ View condition in diffusion model.
$\lambda_{n}$ View condition of $n$-th optimization step.
$r$ Azimuth angle of camera for $\lambda_{n}$.
$\Omega$ Set of views with strong priors, $\Omega=\{\text{front},\text{side}\}$.
$R_{n}$ 2D image rendered from 3D representation at step n1n\!-\!1.
$\mathcal{L}_{\text{P}}$ Similarity-based partial order loss.
$\boldsymbol{z}_{t}$ 2D latent at diffusion step $t$, $t\in\{0,\ldots,T\}$.
$\boldsymbol{z}_{0}$ Clean 2D latent (reverse diffusion output).
$\boldsymbol{z}_{T}$ Final noise latent sampled from $p(\boldsymbol{z}_{T})$.
$T$ Total diffusion time steps.
$p(\cdot)$ Probability density function.
$p_{2D}(\boldsymbol{z}_{0}\mid c)$ 2D latent distribution (conditioned on $c$).
$\tilde{p}_{3D}(\theta)$ Unnormalized PDF of 3D representation.
NN Number of optimisation iterations.
$\theta$ 3D parameters.
$Z_{\theta,n}$ 3D representation at iteration $n$ (parameterized by $\theta$).
$\nabla_{\theta}$ Gradient operator w.r.t. $\theta$.
$M$ Compatibility term ($c_{v}$ vs. $\lambda_{n}$, Eq. (11)).
$\langle\cdot,\cdot\rangle$ Inner product of two vectors.
$\|\cdot\|$ Euclidean norm of a vector.
$|\cdot|$ Absolute value of a scalar.
$\perp$ Orthogonality (e.g., $c_{sbj}\perp c_{v}$).
$\mathbf{V}^{\emptyset}_{\text{key}}$ Keyword embedding (no view description).
$\mathbf{V}^{\text{view}}_{\text{key}}$ Keyword embedding (with specific view description).
$\delta^{\text{view}}_{\text{key}}$ View-specific residual feature.
$Proj_{v}(u)$ Projection of $u$ onto $v$: $Proj_{v}(u)=\frac{u\cdot v}{\|v\|^{2}}v$.
To tackle the multi-face Janus problem, we propose ConsDreamer, a unified method to enhance view semantic clarity. Fig.4 provides an overview of our approach. Section III-A presents a mathematical analysis of the problem, identifying two problematic terms: the conditional and unconditional terms, both of which introduce prior view preferences during the score distillation process. Based on this analysis, Section III-B introduces VDM to eliminate prior view preferences and integrate target view control in the conditional term. Meanwhile, for the unconditional term, Section III-C explores similarity distribution patterns across views in rendered images and proposes a similarity-based partial order loss $\mathcal{L}_{\text{P}}$ to enhance spatial consistency in 3D content. For clarity, the main symbols and notations used in this section are summarised in Table I.
@@ III-A Analysis of the multi-face Janus problem
where $R_{n}$ is a 2D rendered image from the 3D content in the ${(n-1)}^{\text{th}}$ iteration, and $Z_{\theta,n}$ denotes the 3D representation at the $n^{\text{th}}$ iteration. In Eq.(4), the user prompt $c$ is expected only to provide subject information without views information. However, the current pre-trained T2I models incorporate view prior knowledge when “understanding” user prompt $c$ even if $c$ lacks explicit view information. This prior knowledge arises from the prior preference views of training data. Therefore, we decompose $c$ into a subject component $c_{sbj}$ and a view prior component $c_{v}$, which are expected to be orthogonal:
$\tilde{p}_{3D}(\theta)=\prod_{n=1}^{N}p\left(Z_{\theta,n}\mid\lambda_{n},c_{sbj},c_{v},R_{n}\right),\quad\text{s.t.}\,c_{sbj}\perp c_{v},$ (5)
@@
${\rm\log}\tilde{p}_{3D}(\theta)=\sum_{n=1}^{N}{\rm\log}p\left(Z_{\theta,n}\mid\lambda_{n},c_{sbj},c_{v},R_{n}\right),$ (6)
the gradient of $\theta$ denoted as $\nabla_{\theta}{\rm\log}\tilde{p}_{3D}(\theta)$, can be expressed as:
$\nabla_{\theta}{\rm\log}\tilde{p}_{3D}(\theta)=\sum_{n=1}^{N}\nabla_{\theta}{\rm\log}p\left(Z_{\theta,n}\mid\lambda_{n},c_{sbj},c_{v},R_{n}\right)$ (7)
$=\sum_{n=1}^{N}\left(\frac{\partial{\rm\log}p\left(Z_{\theta,n}\mid\lambda_{n},c_{sbj},c_{v},R_{n}\right)}{\partial_{Z_{\theta,n}}}\right)\frac{\partial_{Z_{\theta,n}}}{\partial_{\theta}},$
as $c_{sbj}$ and $c_{v}$ are independent, Eq.(7) can be further expanded using Bayes’ theorem. Specifically, the joint distribution $p(Z_{\theta,n},\lambda_{n},c_{sbj},c_{v},R_{n})$ satisfies the decomposition property, allowing it to be expressed as $p(Z_{\theta,n})p(\lambda_{n},c_{sbj},c_{v},R_{n}\mid Z_{\theta,n})$. This ensures that the transformation using Bayes’ theorem is valid:
$\begin{aligned} \nabla_{\theta}{\rm\log}\tilde{p}_{3D}(\theta)&=\sum\limits_{n=1}^{N}\!\left(\!\frac{\partial{\rm\log}p\left(\lambda_{n},c_{sbj},c_{v},R_{n}\!\mid\!Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}\right.\\ &\quad\left.+\frac{\partial{\rm\log}p\left(Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}\!-\!\frac{\partial{\rm\log}p\left(\lambda_{n},c_{sbj},c_{v},R_{n}\right)}{\partial_{Z_{\theta,n}}}\right)\frac{\partial_{Z_{\theta,n}}}{\partial_{\theta}},\end{aligned}$
the term $\frac{\partial{\rm\log}p\left(\lambda_{n},c_{sbj},c_{v},R_{n}\right)}{\partial_{Z_{\theta,n}}}$ is a constant term independent of $Z_{\theta,n}$, and its derivative is zero, thus can be removed:
$\begin{aligned} \nabla_{\theta}{\rm\log}\tilde{p}_{3D}(\theta)\!=\!\sum\limits_{n=1}^{N}\!\left(\!\frac{\partial{\rm\log}p\left(Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}+\!\frac{\partial{\rm\log}p\left(\lambda_{n},c_{sbj},c_{v},R_{n}\!\mid\!Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}\!\right)\!\frac{\partial_{Z_{\theta,n}}}{\partial_{\theta}},\end{aligned}$
where $\frac{\partial{\rm\log}p\left(Z_{\theta,n}\right)}{\partial_{Z_{\theta,n}}}$ is the unconditional score [ho2020denoising-2405-7], [song2020score-2405-35] modelled by T2I models, intuitively reflects the model’s perception of 3D objects without external view guidance. However, as previously mentioned, such perception is prone to prior preference views, which leads to the multi-face Janus problem. To address this issue, Section III-C explores similarity distribution patterns across rendered images and proposes a similarity-based partial order loss $\mathcal{L}_{\text{P}}$ to enhance the model’s view awareness in the absence of conditional view information.
@@
$\begin{aligned} &\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}p\left(\lambda_{n},c_{sbj},c_{v},R_{n}\mid Z_{\theta,n}\right)\\ &\!=\!\underbrace{\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}p\!\left(c_{sbj},c_{v},R_{n}\!\mid\!Z_{\theta,n}\right)}_{\text{\normalsize Prior View Bias}}\!+\!\underbrace{\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}p\!\left(\lambda_{n},c_{sbj},R_{n}\!\mid\!Z_{\theta,n}\right)}_{\text{\normalsize Target View Control}}\!+\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}M,\end{aligned}$
where $\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}p\!\left(c_{sbj},c_{v},R_{n}\!\mid\!Z_{\theta,n}\right)$ is independent of the view condition $\lambda_{n}$ and denotes the prior view bias term, while $\frac{\partial}{\partial_{Z_{\theta,n}}}{\rm\log}p\!\left(\lambda_{n},c_{sbj},R_{n}\!\mid\!Z_{\theta,n}\right)$ encodes view information exclusively governed by the view condition $\lambda_{n}$, corresponding to the target view control term, and $M=~\frac{p\left(\lambda_{n},c_{sbj},c_{v},R_{n}\mid Z_{\theta,n}\right)}{p\left(c_{sbj},c_{v},R_{n}\mid Z_{\theta,n}\right)p\left(\lambda_{n},c_{sbj},R_{n}\mid Z_{\theta,n}\right)}$.
$c_{sbj}$ and $R_{n}$ are constants in the currentoptimisationn step, then $M$ can be simplified by using the definition of conditional probability:
@@
Observing Eq.(10) and Eq.(11), when the view control $\lambda_{n}$ and the prior view preference $c_{v}$ conflict under the condition $Z_{\theta,n}$, i.e., $p\left(\lambda_{n}\mid c_{v},Z_{\theta,n}\right)\ll p\left(\lambda_{n}\mid Z_{\theta,n}\right)$, the prior view bias and the target view control in Eq.(10) will have a detrimental impact on the 3D generation jointly, which echoes the abnormal activations observed in the subject-keyword CA map in Fig. 3(a). Additionally, in this case, $M$ approaches 0, and the third term $\nabla_{Z_{\theta,n}}{\rm\log}M$ in Eq.(10) introduces a large negative gradient, disrupting the optimisation process of the model. Conversely, when $\lambda_{n}$ and $c_{v}$ are consistent, i.e., $p\left(\lambda_{n}\mid c_{v},Z_{\theta,n}\right)\approx p\left(\lambda_{n}\mid Z_{\theta,n}\right)$, the gradient directions of the first two terms in Eq.(10) align, leading to a more stable optimization process. Furthermore, as $M$ approaches 1, the gradient term $\nabla_{Z_{\theta,n}}{\rm\log}M$ approaches 0, ensuring it does not interfere with the optimisation process. Therefore, establishing semantic consistency between $c_{v}$ and $\lambda_{n}$ is essential for effectively mitigating the multi-face Janus problem. Section III-B eliminates prior view preferences by disentangling view features and injects precise view control when optimising the rendering results for specific views, thereby enhancing the semantic consistency between $c_{v}$ and $\lambda_{n}$.
@@ III-B View Disentanglement Module
Based on the mathematical analysis in Section III-A, we aim to eliminate the prior view preference and strengthen the target view specified by the camera parameters. The optimisation objective can be expressed as maximising the consistency between $c_{v}$ and $\lambda_{n}$:
@@
The process is illustrated in the upper of Fig.4. First, identify the keyword from the prompt (e.g., the keyword “squirrel” of the prompt “A squirrel eating a burger”), and extract the keyword embedding $\mathbf{V}_{{key}}^{\emptyset}$ from the prompt encoding results. Subsequently, new embeddings $\mathbf{V}_{{key}}^{{view}}$ corresponding to various views are obtained by combining the prompt with different view descriptions (e.g., “back view,” “side view,” and “front view”). Unlike $\mathbf{V}_{{key}}^{\emptyset}$, $\mathbf{V}_{{key}}^{{view}}$ incorporates auxiliary contextual information, which enables the view-specific features $\delta_{{key}}^{{view}}$ isolating content-agnostic information. The process is formulated as follows:
$\delta_{{key}}^{{view}}=\mathbf{V}_{{key}}^{{view}}-{Proj}_{\mathbf{V}_{key}^{\emptyset}}\left(\mathbf{V}_{{key}}^{{view}}\right),$ (13)
[…]
@@
$\mathbf{V}^{\emptyset}_{\mathit{key}}\!\leftarrow\!\mathbf{V}^{\emptyset}_{\mathit{key}}\!-\!\sum^{\Omega}\!Proj_{\delta_{{key}}^{{view}}}(\mathbf{V}^{\emptyset}_{\mathit{key}})\!+\!w_{1}\!\cdot\!\frac{|r|}{90}\!\cdot\!\delta^{\mathit{side}}_{\mathit{key}},$ (14)
@@
where $w_{1}$ controls the intensity of view injection, $|r|/90$ normalizes the azimuth $r$ symmetrically within the range of -90° to 90°, ensuring that the injection intensity corresponds proportionally to the deviation from the front view. When the azimuth $r$ falls within the range $(-180^{\circ},-90^{\circ})\cup(90^{\circ},180^{\circ})$, both $\delta_{{key}}^{{back}}$ and $\delta_{{key}}^{{side}}$ must be considered simultaneously, because diffusion models may be biased toward generating front views [armandpour2023re-perpneg]:
$\mathbf{V}^{\emptyset}_{\mathit{key}}\!\leftarrow$ $\mathbf{V}^{\emptyset}_{\mathit{key}}\!-\!\sum^{\Omega}Proj_{\delta_{{key}}^{{view}}}(\mathbf{V}^{\emptyset}_{\mathit{key}})$ (15)
$\!+\!w_{2}\!\cdot\!\frac{|r|\!-\!90}{90}\!\cdot\!\delta^{\mathit{back}}_{\mathit{key}}\!+\!w_{3}\cdot\frac{180\!-\!|r|}{90}\!\cdot\!\delta^{\mathit{side}}_{\mathit{key}},$
where $w_{2},w_{3}$ control the intensity of view injection. As illustrated in Fig.4, we implement the dynamic adjustment of injection intensity using Eq.(14) and Eq.(15), the model can more accurately “understand” the target view $\lambda_{n}$ specified by the rendering camera. We fix $w_{1}=1.0$, $w_{2}=1.5$ and $w_{3}=1.0$ in all experiments, and results in Section IV demonstrate that this approach effectively mitigates the multi-face Janus problem.
Figure 6: Partial order loss for cross-view consistency. The random camera views are distributed on the unit circle based on their azimuth angles. A reference camera (e.g., camera 0) is selected, and a red dotted line parallel to the horizontal axis is constructed from this point to mirror the reference to a symmetric position. The azimuthal distances between other camera views and the nearest reference are computed to determine the expected similarity relationships ($\text{sim}_{0,3}>\text{sim}_{0,2}>\text{sim}_{0,1}$). The right-hand side of the figure computes the actual similarity relationships using cosine similarity, and $\mathcal{L}_{\text{P}}$ evaluates the alignment between the actual and expected similarity relationships.
$\begin{aligned} \mathbf{V}_{{key}}^{prompt}\!\leftarrow\!\mathbf{V}_{{key}}^{prompt}\underbrace{\!-\!\sum^{\Omega}Proj_{\delta_{{key}}^{{view}}}\!\left(\!\mathbf{V}_{{key}}^{prompt}\!\right)\!}_{\text{\scriptsize Prior Elimination}}\underbrace{+\eta\cdot\delta_{{key}}^{{back}}}_{\text{\scriptsize Control Injection}},\end{aligned}$
Figure 7: Impact of $\mathcal{L}_{\text{P}}$ on 3D content over iterations. Using “A squirrel eating a burger” as an example, $\mathcal{L}_{\text{P}}$ demonstrates a significant suppression effect on the multi-face Janus problem encountered during early training stages.
III-C Partial Order Loss for Cross-View Consistency
Figure 8: Qualitative ablation study results (prompt: “A photo of a Pikachu, yellow, electric”). Incrementally integrating our proposed method’s components into the baseline enhances generation consistency. To enhance the persuasiveness of our analysis, we include Perp-Neg as a variable.
To achieve this, we explore the relationship between the multi-view rendered images. First, each rendered image is encoded into high-dimensional semantic features, which capture the global view-aware semantics of the images and exhibit some robustness to flipping, rotation, and symmetry variations of the rendered content [bengio2014representationlearningreviewnew]. A random image is selected from all rendered images as the reference image, and the cosine similarity between the other images and the reference image is computed at the feature level. As shown in Fig.6, the red camera represents the reference view selected from a series of randomly generated camera views $\Lambda$. It is evident that the closer the azimuth is to the reference azimuth, the higher the similarity score between the corresponding rendered image and the reference image. Furthermore, the scores gradually decrease symmetrically as the azimuth angle deviates from the reference. Therefore, we define a similarity-based partial order loss, where the specific ordering method is illustrated in Fig.6. A Cartesian coordinate system is established, and the random camera views are distributed on the unit circle based on their azimuth angles. One camera view is randomly chosen as the reference view, e.g., camera 0 in Fig.6, and a red dotted line parallel to the horizontal axis is constructed from this point. The reference camera is then mirrored to a symmetric point using this line. The azimuthal distances between other camera views and the nearest reference camera view are computed, determining the expected similarity partial order among the corresponding rendered images: $\text{sim}_{0,3}>\text{sim}_{0,2}>\text{sim}_{0,1}$. The similarity relationship between the features of rendered images obtained from Eq.(9) must also conform to this order. Accordingly, the cosine similarity between the semantic features of corresponding rendered images is computed as the actual similarity and a similarity-based partial order loss $\mathcal{L}_{\text{P}}$ is employed to constrain the actual similarity to align with the expected similarity relationships. $\mathcal{L}_{\text{P}}$ is defined as:
$\mathcal{L}_{\text{P}}\!=\!\sum_{i=1}^{|\Lambda|-2}\!\max\!\left(0,\text{sim}\!\left(F^{i+1},F^{0}\right)\!-\!\text{sim}\!\left(F^{i},F^{0}\right)\right),$ (17)
TABLE II: Quantitative Results of Ablation Study